Order Ancef 200 mg IV q 4h for a child weighing 22 lbs. The recommended dose...

  1. Order Ancef 200 mg IV q 4h for a child weighing 22 lbs.
    1. The recommended dose is 80 – 160 mg / kg / day
    2. Is the dose safe?
    3. Ancef is provided as 1 g in 10 mL
    4. How many mL would you draw up?
    5. You are to add the Ancef to 10 mL of IV fluid and infuse over ½ hour.
    6. How fast would you set the IVAC at?
Answer #1

ANSWER.

b) Calculation of dose for this patient is safe or not.

          Ordered dose = 200 mg IV (4h).

          Weight of the patient = 22 lb.

     Convert lb to kg for calculation

                 1 lb = 0.454 kg.

           Weight of the patient = 22 * 0.454

                                              = 10kg.

      Recommended dose = 80-160 mg/kg/day.

       Dose required for this patient = 80 * 10 - 160 * 10

                                                       = 800 - 1600 mg/day.

       Dose received by the patient as per ordered dose = 200 * 6 (every 4 hr)

                                                                                       = 1200 mg/day.

        * Ordered dose(200mg q4h) is safe for this patient.

d) Calculation of volume(ml) of drug required .

          Formula = Desired dose/Dose in hand * Quantity.

     Here

         Desired dose(D) = 200 mg.

         Dose in hand (H) = 1000 mg (1g).

           Quantity (Q) = 10 ml.

   As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 200/1000 * 10

                                                        = 0.2 * 10

                                                        = 2 ml.

f) Calculation of rate(ml/hr) set in the IVAC pump.

        Formula

          Rate(ml/hr) = Volume(ml) /Time(hr).

     Here

            Volume(ml) = 10 ml.

             Time (hr) = 0.5 hr (30 min).

   As per above formula

            Rate (ml/hr) = 10 / 0.5

                                = 20 ml/hr.